OpenStudy (kamekona):

What is square root of 114 approximated to the nearest tenth?

OpenStudy (boldjon):

\[\sqrt{114}\approx10.7\]

jhonyy9 (jhonyy9):

so you know that 10*10=100 and 11*11=121 - how you wan round it down so than 10 or up and than will be 11 ok.?

OpenStudy (kamekona):

thanks

jhonyy9 (jhonyy9):

was my pleasure and anytime good luck bye